OpenStudy (anonymous):

Which of these sets of numbers contain no irrational #'s a.-5 squared -196 squared -8.15 b.-144 squared 49 squared 6.6 c.-2, 3/8, 10 squared, -0.6868 D.-1, 5/6, 11 squared

OpenStudy (anonymous):

pLEASE HELP!!!!!

OpenStudy (anonymous):

sure

OpenStudy (paxpolaris):

ist that wat option a. looks like?? are there no square roots ( \(\sqrt{\ }\) ) ?? a. \(-5^2,-196^2,-8.15\)

OpenStudy (anonymous):

C duh...

OpenStudy (anonymous):

no if it says squared next to it then it is a squared root

OpenStudy (paxpolaris):

why c????

OpenStudy (anonymous):

C and D actually

OpenStudy (paxpolaris):

can u attach picture

OpenStudy (anonymous):

is the -0.6868 reoccuring?

OpenStudy (anonymous):

or exactly?

OpenStudy (anonymous):

ok I will show u

OpenStudy (anonymous):

if it's reoccuring, then your answer is D my friend

OpenStudy (anonymous):

thanks

OpenStudy (paxpolaris):

i can't see any irrational numbers the way its written?????

OpenStudy (paxpolaris):

sorry just got it, my answer would be: b. \(\large-\sqrt{144},\sqrt{49},6.6\) because it equals... -12, 7, 6.6

OpenStudy (paxpolaris):

in a. \(\sqrt5\) is irrational in c. \(\sqrt{10}\) in d. \(\sqrt{11}\) only b has no irrationals....